OpenStudy (anonymous):

how do you solve this problem 3/x-2+4/5=27/5(x-2)

OpenStudy (anonymous):

\[\frac{ 3 }{ x-2 }+\frac{ 4 }{ 5}=\frac{ 27 }{ 5(x-2) }\] is this your problem? if so you need to find the lcd of all 3 fractions let me know what you come up with and we will see what corrections need to be made